# Zero-sum partitions of Abelian groups and their applications to magic-type labelings

### v3

@@ 4. Final Remarks
Finally, let us complement Definition 1.4. Let $S$ be a subset of cardinality $m=6k+s$, for a positive integer $k$ and $s\in\{0,2,4\}$, of a finite Abelian group $\Gamma$. A partition of $S$ into $k$ good $6$-subsets and $s/2$ zero-sum $2$-subsets is called a Skolem partition of $S$. It is known that $\mathbb{Z}_{m}$, for $m=6k+s$ with any positive integer $k$, has a Skolem partition when $s\in\{0,4\}$, but the situation is slightly different for $s=2$. For $m\equiv 2$ or $8\pmod{24}$, there exists a Skolem partition, whereas for $m\equiv 14$ or $20\pmod{24}$, such a partition does not exist (see Lemma 4 in [22]). Note that there exist groups $\Gamma$ with $|I(\Gamma)|>1$ such that the set $R=\Gamma^{*}\setminus I(\Gamma)$ has a Skolem partition. For instance, let $\Gamma\cong(\mathbb{Z}_{2})^{\eta}\times H$ for some natural number $\eta>1$ and $|H|\equiv 1\pmod{6}$. By Theorem 1.5, there exists a partition of $H^{*}$ into good $6$-subsets:
@@
As reflected in Proposition 3.7, given a graph $G=(V,E)$ with a partition of the vertex set into subsets of false twins, existence of a zero-sum partition of an Abelian group $\Gamma$ into subsets of corresponding cardinalities is a sufficient condition for existence of a $\Gamma^{*}$-distance magic labeling of $G$. One can easily see that this condition is not necessary. For example, consider the group $\mathbb{Z}_{6}$. There exists no zero-sum partition of $\mathbb{Z}_{6}^{*}$ as required in Proposition 3.7. On the other hand, $\{\{1,2\},\{3\},\{4,5\}\}$ is a constant-sum partition of $\mathbb{Z}^{*}_{6}$, with constant sum $3$, which leads to a $\mathbb{Z}_{6}^{*}$-distance magic labeling of the complete tripartite graph $K_{1,2,2}$. So far, constant-sum partitions of Abelian groups have been studied only for the case of partitioning into three subsets (see [5]). It would be interesting to consider constant-sum partitions into more than three subsets.

### v4

@@ 4. Final Remarks
Finally, let us complement Definition 1.4. Let $S$ be a subset of cardinality $m=6k+s$, for a positive integer $k$ and $s\in\{0,2,4\}$, of a finite Abelian group $\Gamma$. A partition of $S$ into $k$ good $6$-subsets and $s/2$ zero-sum $2$-subsets is called a Skolem partition of $S$. It is known that $\mathbb{Z}_{m}^{*}\setminus I(\mathbb{Z}_{m})$, for $m=6k+s$ with any positive integer $k$, has a Skolem partition when $s\in\{0,4\}$, but the situation is slightly different for $s=2$. For $m\equiv 2$ or $8\pmod{24}$, there exists a Skolem partition, whereas for $m\equiv 14$ or $20\pmod{24}$, such a partition does not exist (see Lemma 4 in [22]). Note that there exist groups $\Gamma$ with $|I(\Gamma)|>1$ such that the set $R=\Gamma^{*}\setminus I(\Gamma)$ has a Skolem partition. For instance, let $\Gamma\cong(\mathbb{Z}_{2})^{\eta}\times H$ for some natural number $\eta>1$ and $|H|\equiv 1\pmod{6}$. By Theorem 1.5, there exists a partition of $H^{*}$ into good $6$-subsets:
@@
As reflected in Proposition 3.8, given a graph $G=(V,E)$ with a partition of the vertex set into subsets of false twins, existence of a zero-sum partition of an Abelian group $\Gamma$ into subsets of corresponding cardinalities is a sufficient condition for existence of a $\Gamma^{*}$-distance-magic labeling of $G$. One can easily see that this condition is not necessary. For example, consider the group $\mathbb{Z}_{6}$. There exists no zero-sum partition of $\mathbb{Z}_{6}^{*}$ as required in Proposition 3.8. On the other hand, $\{\{1,2\},\{3\},\{4,5\}\}$ is a constant-sum partition of $\mathbb{Z}^{*}_{6}$, with constant sum $3$, which leads to a $\mathbb{Z}_{6}^{*}$-distance-magic labeling of the complete tripartite graph $K_{1,2,2}$. So far, constant-sum partitions of Abelian groups have been studied only for the case of partitioning into three subsets (see [5]). It would be interesting to consider constant-sum partitions into more than three subsets.